# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v4

@@ 3. Unified Space
We proceed to develop a methodology that integrates code transformation, tile size selection and pragma insertion into a unified optimization framework. This involves implementing maximal distribution for code transformation, followed by the design and implementation of a template capable of executing code transformation, tiles size selection and pragma insertion simultaneously. We illustrate our code transformation process step by step using the example from Listings 1 to 2. Input  Sisyphus takes an affine C/C++ file as input and automatically extracts a polyhedral representation of the code using PoCC (PoCC, [n. d.]). This representation allows us to gather all relevant information from the code, such as loop order, loop bounds, array sizes, and more. We will use this information to create the template and to automatically generate the NLP discussed in Section 4. Our framework will be particularly effective for codes that are fully distributable and fully permutable. However, in cases where the transformations described later are not legally possible, our framework still functions but restricts the solution space more. Maximal Distribution  To kickstart the process, we begin by thoroughly distributing the program, with the goal of maximizing the distance between each statement. This distribution strategy creates ample opportunities for parallelization by maximizing the distance between dependencies (reverse of fusion). Typically, this results in one statement per loop body with perfectly nested loops. To achieve this, we employ ISCC (Verdoolaege, 2011). It allows us to explore various schedule distributions and validate the legality of transformations by ensuring dependency constraints are preserved. The code presented in Listing LABEL:lst:distributed displays the fully distributed version of the code from Listing 1. Strip-mining  After maximal distribution, we apply strip-mining to each loop twice. This process transforms the original loop, such as the loop L0 in Listing LABEL:lst:distributed, into three loops with trip counts $I0\_S0$, $I1\_S0$, and $I2\_S0$ (L0, L1, L2 in Listing LABEL:lst:strip). The product of these three loop trip counts equals the original trip count, i.e., $I0\_S0\times I1\_S0\times I2\_S0=200$. Since strip mining is inherently legal, there is no necessity to validate the legality of this transformation. Subsequently, we can arrange the loops in different permutations if it is legal. After applying three levels of strip-mining, the code from Listing LABEL:lst:distributed is transformed into the code seen in Listing LABEL:lst:strip. Creation of different levels  Next, we consider possible permutations of the strip-mined loops. If these permutations are legal, we establish three loop levels. We ensure legality by checking dependency preservation using ISCC (Verdoolaege, 2011). Level 2:  The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize a tree reduction if a reduction loop is unrolled and the option is enabled. For example, this is illustrated by loops L9, L10, and L11 in Listing 2. Level 1:  The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations. In cases where loop permutations allow us to achieve perfectly nested loops suitable for pipelining, we restrict pipelining to a single loop to ensure efficient synthesis. Flattening the loops at Level 1 into a single loop and then pipelining the resulting loop introduces significant design complexity and leads to excessively long synthesis times. Thus, if a loop at Level 1 is pipelined, it implies that the trip counts of the other loops at the same level are reduced to one, ensuring manageable design complexity and synthesis times. For instance, in Listing 2, the loop $j$ (L8) is pipelined for the second statement (S1). This results in the loops $k$ and $i$ (L10 and L11 in Listing 3 ) having a trip count of 1, enabling them to be eliminated from the code. Level 0:  Finally, the outermost level manages coarse-grained unrolling, tiling, and/or sequential execution. It not only controls parallelism at a coarse level but also determines the on-chip buffer size through tiling. As described in the following paragraph, the loop order at this level is included in the design space. This level is illustrated by loops L5, L6, and L7 in Listing LABEL:lst:small_mem. The loop bounds determine the on-chip buffer sizes for arrays A and B, transferred below loop L5. Resource constraints also influence the sequential execution of loop L7. If the loop body contains a single statement and loop permutation is feasible, we disable coarse-grained unrolling. This achieves equivalent results since our template allows fine-grained unrolling of the innermost loop. Loop Permutation  Loop order selection is not required for levels 1 and 2, as only a single loop is pipelined, and the innermost level (level 2) is fully unrolled, making the loop order irrelevant. Conversely, at the outermost level (level 0), all permutations are considered, and the NLP selects the permutation. In Listings LABEL:lst:small_mem and 2, we can see that the NLP selected two different permutations for the statement S1 (loops L5, L6, and L7).
@@ 20 C[i][j]*=beta; // S0
Listing 3: Template
Fusion  We have opted not to incorporate fusion into our model for several reasons. Firstly, our objective is to maximize parallelization for computation-bound kernels. Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities. Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation. Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times. Template  In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations. The iterator of the loop retains its original name, with an added number indicating the level. For example, loop L14 in Listing 3 refers to loop L2 in Listing 1 at level 2 of the strip mining. Notably, our framework is capable of preserving the original loop order, allowing it to retain the existing loop structure if the code is already optimized. Space  Therefore, the challenge involves determining the trip counts of each loop (e.g., $J0\_S0$) and the loop order of the level 0. Additionally, the design must adhere to resource constraints, including the usage of DSP and on-chip memory. Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources. The on-chip buffers size must not exceed the on-chip memory capacity. Consequently, for a large problem size, the outermost level will enable control over the buffer sizes.
@@ 4. NLP
Presented here is a comprehensive set of constraints and variables employed in a nonlinear program aimed at discovering the theoretical solution space outlined in Section 3. We employ the methodology proposed by (Pouget et al., 2024b, a), adapted to our specific context. Similar to their approach, we establish an estimation of the latency by assuming optimistic DSP utilization, assuming perfect resource reuse. However, we afford the user the flexibility to adjust the DSP limit or opt for a pessimistic DSP utilization scenario, where no reuse between statements is considered. Moreover, users have the ability to adjust the size of on-chip memory, define the maximum number of array partitions, and customize the latency and resources allocated for each operation. This flexibility ensures adaptability across various platforms and compilers. To gather all the required information for the NLP, we utilize PoCC (PoCC, [n. d.]) to extract the intermediate representation.
@@ 4.1. Variables
Table 2 delineate the sets, variables, and constants utilized in our NLP formulation.
@@ 4.2. Constraints
Now, we explore the precise meaning and implications of each constraint. To make this more understandable, we provide examples using the listings presented in the paper. Trip Count  Equation 1 constrains the trip count of each loop, ensuring that the product of the trip counts equals the original trip count. The three loops after strip-mining, e.g., L6, L7, and L8 in Listing LABEL:lst:strip, must be equivalent to the original loop L2 in our example from Listing LABEL:lst:distributed. Therefore, we require $I0\_S0\times I1\_S0\times I2\_S0=200$.
@@
Pipeline  Constraints 2, 3, 4, and 5 facilitate the selection of loop pipelining. Constraint 2 ensures that each loop has a boolean flag indicating whether the original loop is pipelined. For example, in Listings LABEL:lst:small_mem and 2, the statement $S1$ shows that the loop $j1$ (L8) is pipelined, while the original loop is $j$ (L3 in Listing 1), indicating that $loop_{j}\_pip=1$. As explained in Section 3, only one loop can be pipelined at each middle level. This means the sum of the boolean flags must be less than or equal to 1 (Eq. 4), and the trip count of all other non-pipelined loops at the middle level must be 1 (Eq. 3). For example, in the template (Listing 3), only loop L9 is pipelined at level 1 of S1 (loops L9, L10, and L11), so $loop_{j}\_pip=1$. Consequently, $loop_{i}\_pip=0$ and $loop_{k}\_pip=0$ (Eq. 4), and Eq. 3 indicates that the trip counts of loops L10 and L11 are 1, allowing their removal.
@@
$\mathcal{V}$ Level of the strip mining, 0 for coarse-grained/sequential, 1 for pipeline and 2 for fine-grained unrolled
@@
$loop_{l}\_UF$ Coarse-grained unroll factor of the loop $l$ at level 0 of the strip-mining
$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined at level 1 of the strip-mining
$cache_{l}\_arr_{a}$ Boolean to know if the array $a$ is transferred on-chip after the loop $l$ at level 0 of the strip-mining
@@
Table 2. Overview of the set, constant and variables employed in formulating the NLP
Eq. 5 allows the computation of the initiation interval (II), which is determined based on the loop’s properties and the iteration latency of the reduction operation. Specifically, when the loop being pipelined is a reduction loop, this follows the approach outlined in (Pouget et al., 2024b, a). In Listings LABEL:lst:small_mem and 2, the pipelined loops are non-reduction loops with no dependencies when iterating, hence $II=1$. However, if loop L10 in Listing 3 were pipelined, it would involve a reduction, meaning that at each iteration, C[i][j] needs to be read and written. Thus, to read for iteration $k+1$, we must wait for the write to finish at iteration $k$, which takes $IL_{+}$ cycles (latency of the reduction operation).
@@
Coarse-grained unrolling  We enforce coarse-grained unrolling solely for non-reduction loops (Eq. 8). For example, the loop L3 in Listing LABEL:lst:distributed cannot be coarse-grained unrolled as it is a reduction loop. Additionally, the unroll factor (UF) must divide the trip count of the current loop and be less than or equal to it (Eq. 9 and 6). However, as elaborated in Section 3, coarse-grained unrolling is disregarded if there is only one statement within the loop (Eq. 7). This implies that for Listing 1 coarse-grained unrolling is disabled.
@@
(9) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l,0}==0$
On-chip memory  Constraints 10 and 11 ensure that each array, represented by a boolean (Eq. 10), is cached on-chip at a single location within the same loop body, as only one boolean can be true for each loop body (Eq. 11). In Listing 3, for instance, array $A$ can be cached either before any loops (line 5) or after L6, L7, or L8 (lines 23, 25, and 27). In Listing 2, $A$ is cached on-chip with the on-chip footprint matching the size of the original array (line 2). Meanwhile, in Listing LABEL:lst:small_mem, $A$ is cached on-chip at line 18, but with a smaller on-chip footprint compared to the original. Based on where each array is transferred, we can compute the on-chip footprint. For example, in Listing LABEL:lst:small_mem, array $A$ is transferred below loop L5, which has a trip count of 60, so the footprint of $A$ on-chip is $200\times\frac{240}{60}$.
@@
Array partitioning  We consider unrolling only if all the data needed for unrolling can be accessed in parallel, i.e., these data are on different on-chip banks thanks to the pragma array_partitioning. For each array $a$ and each dimension $d$ of this array, we want the factor of the cyclic array partitioning for this dimension to be greater than the unroll factor (Eq. 15) to ensure that all data are on different banks. If the array is reused, we need to ensure that the array partitioning is optimal for each statement. Therefore, we force the array partitioning to be a multiple of each unroll factor that iterates through this dimension (Eq. 14). $AP_{a,d}$ gives the cyclic array partitioning factor for array $a$ in dimension $d$ that we use in the generated code. Eq. 13 assures that the array partitioning per array (product of the partitioning factors per dimension) is less than or equal to the limit fixed by the user. For example, in Listing 2, $C$ is unrolled 200 times in the first dimension (lines 11, 12) and 4 times in the second dimension (lines 13, 14) for the first loop body, and 200 times in the first dimension (lines 25, 26) for the second loop body. Thus, we have $AP_{C,1}=200$ and $AP_{C,2}=4$, and $200\times 4=800\leq max_{part}$, where $max_{part}$ is 1,024 in our case.
@@
DSP utilization  Finally, we need to manage DSP utilization. Given the challenges in predicting DSP reuse, we use a boolean flag, $reuse_{opt}$, set by the user to determine the level of DSP reuse. If $reuse_{opt}=1$, we assume an optimistic reuse scenario: DSPs used in one loop body can be directly reused by subsequent loop bodies if the operations are identical. However, if the operations differ, as discussed in (Pouget et al., 2024b, a), no reuse is assumed. Therefore, the total DSP usage is calculated as the sum of the maximum DSPs required for each operation across all loop bodies (Eq. 16). If $reuse_{opt}=0$, we assume no reuse between different loop bodies, and we sum the DSP requirements for each statement (Eq. 17). Based on the value of $reuse_{opt}$, the total DSP usage must be less than or equal to the available DSPs (Eq. 18 and 19). For instance, if a multiplication operation (*) uses 3 DSPs and an addition operation ($+$) uses 2 DSPs, in Listing 2, S0 requires $200\times 4\times 3$ DSPs for multiplications, and S1 requires $200\times 4\times 3+200\times 4\times 3+200\times 4\times 2$ DSPs for two multiplications and one addition. With optimistic reuse for the same operation (e.g., multiplication), the DSP usage is calculated as: $\max(200\times 4\times 3,2\times 200\times 4\times 3)+200\times 4\times 2$. With pessimistic reuse, the DSP usage is: $200\times 4\times 3+2\times(200\times 4\times 3)+200\times 4\times 2$.
@@ 4.3. Objective Function
We structure the objective function akin to the approach described in (Pouget et al., 2024b, a), tailoring it to meet the specific needs of our problem. Here, $Lat^{s}_{2}$ corresponds to the fine-grained unrolled level (level 2) for the statement $s$, $Lat^{s}_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled level (level 1), and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching (level 0). The memory latency $Lat^{s}_{mem}$ for a statement $s$ corresponds to the time required to transfer the array from off-chip to on-chip and back for the arrays needed by a statement $s$. We assume the memory transfer is pipelined, allowing the transfer of one data element (with a maximum bitwidth of 512 bits) per cycle. Therefore, the latency at a given level is calculated as the array footprint divided by the burst size (bitwidth used for transferring the array), multiplied by the trip count of the loop that iterates over the function responsible for the transfer, based on the selected permutation of level 0. Additionally, if an array is fully transferred on-chip for a statement $s$ and reused by a subsequent statement $s^{’}$, no further transfers are needed. The burst size for an array must be a divisor of its last dimension. For instance, if we have an array $float\text{ }A[512][20]$, the maximum burst size of this array $A$ is 128 bits, since the last dimension, $20\times 32$ bits, can be divided by the largest power of 2, which is 128.
@@
Finally, the objective function (Eq. 20) is defined as the sum of the latencies of each loop body.
(20) $obj\_func=\sum_{s\in\mathcal{S}}Lat_{s}$
5. Post-Optimization and Usage

### v5

@@ 3. Unified Space
We proceed to develop a methodology that integrates code transformation, tile size selection and pragma insertion into a unified optimization framework. This involves implementing maximal distribution for code transformation, followed by the design and implementation of a template capable of executing code transformation, tiles size selection and pragma insertion simultaneously. We illustrate our code transformation process step by step using the example from Listings 1 to 2. Input  Sisyphus takes an affine C/C++ file as input and extracts a polyhedral representation using PoCC (PoCC, [n. d.]), gathering information like loop order, bounds, and array sizes. This data is used to create the template and generate the NLP. The framework is most effective for fully distributable and permutable codes but still works with restricted solution space when transformations are not fully legal. Maximal Distribution  To initiate the process, we distribute the program to maximize the distance between statements, promoting parallelization by reducing dependency overlap. Typically, this yields one statement per loop body within perfectly nested loops. Using ISCC (Verdoolaege, 2011), we explore scheduling options and verify transformation legality by preserving dependencies. Listing LABEL:lst:distributed shows the fully distributed version of the code from Listing 1. Strip-mining  After maximal distribution, we apply strip-mining to each loop twice. This process transforms the original loop, such as the loop L0 in Listing LABEL:lst:distributed, into three loops with trip counts $I0\_S0$, $I1\_S0$, and $I2\_S0$ (L0, L1, L2 in Listing LABEL:lst:strip). The product of these three loop trip counts equals the original trip count, i.e., $I0\_S0\times I1\_S0\times I2\_S0=200$. Since strip mining is inherently legal, there is no necessity to validate the legality of this transformation. Subsequently, we can arrange the loops in different permutations if it is legal. After applying three levels of strip-mining, the code from Listing LABEL:lst:distributed is transformed into the code seen in Listing LABEL:lst:strip. Creation of different levels  Next, we consider possible permutations of the strip-mined loops. If these permutations are legal, we establish three loop levels. We ensure legality by checking dependency preservation using ISCC (Verdoolaege, 2011). Level 2:  The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize a tree reduction if a reduction loop is unrolled and the option is enabled. For example, this is illustrated by loops L9, L10, and L11 in Listing 2. Level 1:  The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations. In cases where loop permutations allow us to achieve perfectly nested loops suitable for pipelining, we restrict pipelining to a single loop to ensure efficient synthesis. Flattening the loops at Level 1 into a single loop and then pipelining the resulting loop introduces significant design complexity and leads to excessively long synthesis times. Thus, if a loop at Level 1 is pipelined, it implies that the trip counts of the other loops at the same level are reduced to one. For instance, in Listing 2, the loop $j$ (L8) is pipelined for the statement S1. This results in the loops $k$ and $i$ (L10 and L11 in Listing 3 ) having a trip count of 1, enabling them to be eliminated from the code. Level 0:  Finally, the outermost level manages coarse-grained unrolling, tiling, and/or sequential execution. It not only controls parallelism at a coarse level but also determines the on-chip buffer size through tiling. As described in the following paragraph, the loop order at this level is included in the design space. This level is illustrated by loops L5, L6, and L7 in Listing LABEL:lst:small_mem. The loop bounds determine the on-chip buffer sizes for arrays A and B, transferred below loop L5. Resource constraints also influence the sequential execution of loop L7. If the loop body contains a single statement and loop permutation is feasible, we disable coarse-grained unrolling. This achieves equivalent results since our template allows fine-grained unrolling of the innermost loop. Loop Permutation  Loop order selection is not required for levels 1 and 2, as only a single loop is pipelined, and the innermost level (level 2) is fully unrolled, making the loop order irrelevant. Conversely, at the outermost level (level 0), all permutations are considered, and the NLP selects the permutation. In Listings LABEL:lst:small_mem and 2, we can see that the NLP selected two different permutations for the statement S1 (loops L5, L6, and L7). Fusion  We have opted not to incorporate fusion into our model for several reasons. Firstly, our objective is to maximize parallelization for computation-bound kernels. Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities. Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation. Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times. Overal Template  In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations. The iterator of the loop retains its original name, with an added number indicating the level. For example, loop L14 in Listing 3 refers to loop L2 in Listing 1 at level 2 of the strip mining. Notably, our framework is capable of preserving the original loop order, allowing it to retain the existing loop structure if the code is already optimized. Design Space  The challenge is to determine loop trip counts (e.g., $J0\_S0$) and level 0 loop order while meeting DSP and on-chip memory constraints. Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources. The on-chip buffers size must not exceed the on-chip memory capacity. Consequently, for a large problem size, the outermost level will enable control over the buffer sizes.
@@ 20 C[i][j]*=beta; // S0
Listing 3: Template automatically generated for the gemm input code (Listing 1)
@@ 4. NLP
We automatically generate constraints and variables for a nonlinear program to explore the theoretical solution space outlined in Section 3, adapting the approach from (Pouget et al., 2025, 2024) to our context. Latency is estimated with optimistic DSP utilization, assuming perfect resource reuse, but users can adjust the DSP limit or choose a pessimistic scenario with no reuse. Users can also customize on-chip memory size, maximal array partitioning, and latency/resources per operation, ensuring adaptability to different platforms and compilers. To gather all the required information for the NLP, we utilize PoCC (PoCC, [n. d.]) to extract the intermediate representation.
@@ 4.1. Variables
Table 2 defines the sets, variables, and constants utilized in our NLP formulation.
@@ 4.2. Constraints
We describe each constraint’s meaning and implications with examples from the paper’s listings. Trip Count  Equation 1 constrains the trip count of each loop, ensuring that the product of the trip counts equals the original trip count. The three loops after strip-mining, e.g., L6, L7, and L8 in Listing LABEL:lst:strip, must be equivalent to the original loop L2 in our example from Listing LABEL:lst:distributed. Therefore, we require $I0\_S0\times I1\_S0\times I2\_S0=200$.
@@
$\mathcal{V}$ Strip mining level: 0 for coarse-grained/sequential, 1 for pipeline, 2 for fine-grained unrolled
@@
$loop_{l}\_UF$ Coarse-grained unroll factor of the loop $l$ at level 0
$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined at level 1
$cache_{l}\_arr_{a}$ Boolean to know if the array $a$ is transferred on-chip after the loop $l$ at level 0
@@
Table 2. Overview of the sets, constants and variables employed in formulating the NLP
Pipeline  Constraints 25 facilitate the selection of loop pipelining. Constraint 2 ensures each loop has a flag indicating if it is pipelined. For example, in Listings LABEL:lst:small_mem and 2, $S1$ shows loop $j1$ (L8) is pipelined, while its original loop is $j$ (L3 in Listing 1), making $loop_{j}\_pip=1$. As per Section 3, only one loop at level 1 can be pipelined, requiring the sum of pipeline flags must be less than or equal to 1 (Eq. 4). Non-pipelined loops must have a trip count of 1 (Eq. 3). For instance, in Listing 3, only loop L9 is pipelined at level 1 of $S1$, so $loop_{j}\_pip=1$, while $loop_{i}\_pip=0$ and $loop_{k}\_pip=0$. Loops L10 and L11 have trip counts of 1 and can be removed. Eq. 5 calculates the initiation interval (II) based on loop properties and reduction operation latency. In Listings LABEL:lst:small_mem and 2, pipelined loops are non-reduction loops with no dependencies, giving $II=1$. However, if loop L10 in Listing 3 were pipelined, it would involve a reduction requiring $II\geq IL_{+}$, as the write of iteration $k$ must complete before the read of $k+1$.
@@
Coarse-grained unrolling  Coarse-grained unrolling applies only to non-reduction loops (Eq. 6). For instance, loop L3 in Listing LABEL:lst:distributed cannot be unrolled as it is a reduction loop. The unroll factor (UF) must divide the loop trip count and not exceed it (Eq. 7 and 8). However, as noted in Section 3, coarse-grained unrolling is skipped if the loop has only one statement (Eq. 9), as in Listing 1.
(7) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l,0}==0$
@@
On-chip memory  Constraints 10 and 11 ensure each array, represented by a boolean (Eq. 10), is cached at a single location per loop body, with only one boolean true per loop (Eq. 11). In Listing 3, array $A$ can be cached before loops (line 5) or after L6, L7, or L8 (lines 23, 25, 27). In Listing 2, $A$ is cached on-chip, matching the original array’s size (line 2). In Listing LABEL:lst:small_mem, $A$ is cached on-chip at line 18 with a smaller footprint. The on-chip footprint depends on the transfer location. For Listing LABEL:lst:small_mem, $A$ is transferred below loop L5 (trip count 48), so its footprint is $200\times\frac{240}{48}$.
@@
Array partitioning  Unrolling is applied only if all required data can be accessed in parallel, ensured by pragma array_partitioning. For each array $a$ and dimension $d$, the cyclic partitioning factor must exceed the unroll factor (Eq. 13) to ensure data is on separate banks. If the array is reused, partitioning must align with each unroll factor iterating through the dimension (Eq. 14). $AP_{a,d}$ represents the cyclic partitioning factor used in the code, with the total partitioning per array constrained by the user-defined limit (Eq. 15). For example, in Listing 2, $C$ is unrolled 200 times in the first dimension (lines 11, 12) and 4 times in the second (lines 13, 14) for the first loop body, and 200 times in the first dimension (lines 25, 26) for the second. This gives $AP_{C,1}=200$, $AP_{C,2}=4$, and $200\times 4=800\leq max_{part}$, with $max_{part}=1,024$.
@@
DSP utilization  To manage DSP utilization, we use a user-defined flag, $reuse_{opt}$. If $reuse_{opt}=1$, optimistic reuse is assumed: identical operations reuse DSPs across loop bodies, with total usage given by the maximum DSPs per operation (Eq. 16). However, if the operations differ, as discussed in (Pouget et al., 2025, 2024), no reuse is assumed. If $reuse_{opt}=0$, no reuse is assumed, and DSPs are summed for each statement (Eq. 17). In both cases, total usage must not exceed available DSPs (Eq. 18 and 19). If a multiplication (*) uses 3 DSPs and an addition ($+$) uses 2 DSPs, in Listing 2, S0 needs $200\times 4\times 3$ DSPs for multiplications, while S1 requires $200\times 4\times 3+200\times 4\times 3+200\times 4\times 2$ DSPs for two multiplications and one addition. With optimistic reuse, DSP usage for multiplication is $\max(200\times 4\times 3,2\times 200\times 4\times 3)+200\times 4\times 2$, while pessimistic reuse results in $200\times 4\times 3+2\times(200\times 4\times 3)+200\times 4\times 2$.
@@ 4.3. Objective Function
We use the objective function similar to the approach described in (Pouget et al., 2025, 2024), tailoring it to meet the specific needs of our problem. Here, $Lat^{s}_{2}$ corresponds to the fine-grained unrolled level (level 2) for the statement $s$, $Lat^{s}_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled level (level 1), and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching (level 0). The memory latency $Lat^{s}_{mem}$ for a statement $s$ corresponds to the time required to transfer the array from off-chip to on-chip and back for the arrays needed by a statement $s$. We assume the memory transfer is pipelined, allowing the transfer of one data element (with a maximum bitwidth of 512 bits) per cycle. Therefore, the latency at a given level is calculated as the array footprint divided by the burst size (bitwidth used for transferring the array), multiplied by the trip count of the loop that iterates over the function responsible for the transfer, based on the selected permutation of level 0. Additionally, if an array is fully transferred on-chip for a statement $s$ and reused by a subsequent statement $s{’}$ ’, no further transfers are needed. The burst size for an array must divide its last dimension. For instance, if we have an array $float\text{ }A[512][20]$, the maximum burst size for this array is $128$ bits. This is because $128$ is the largest power of $2$ that can evenly divide the size of the last dimension ($20\times 32$ bits).
@@
Finally, the objective function $obj\_func=\sum_{s\in\mathcal{S}}Lat_{s}$ is defined as the sum of the latencies of each loop body.
5. Post-Optimization and Customization